OpenStudy (electronia):

whats 18/30s common denominator.

OpenStudy (electronia):

Medal and fan.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

I need more information.

OpenStudy (anonymous):

2 goes into 18 and 30

OpenStudy (anonymous):

There needs to be two fractions to find a common denominator.

OpenStudy (alexandervonhumboldt2):

hint: 18=2*3*3 30=2*3*5

OpenStudy (mathmale):

I think you mean "greatest common factor of the numerator and denominator of 18/30" not common denominator. Try again. "Reduce" 18/30 to lowest form.

OpenStudy (anonymous):

ARE YOU TRYING TO FIND THE LEAST COMMON DENOMINATOR OR THE MOST COMMON DENOMNATOR?

OpenStudy (mathmale):

Neither. Find the greatest common FACTOR.

OpenStudy (alexandervonhumboldt2):

\[\frac{ 18 }{ 30 }=\frac{ \cancel {2*3}*3 }{ \cancel {2*3}*5 }\]

OpenStudy (mathmale):

Nice work, Alex! Thanks for showing how you arrived at that result. Now, @Electronia , please do the same thing to each of the given possible answers. If your results include 3/5, you've found a solution; otherwise y ou have not.

OpenStudy (mathmale):

Please show your work.

OpenStudy (alexandervonhumboldt2):

thx mathmale :)

OpenStudy (electronia):

I can, I need a answer BIGGER than 18/30

OpenStudy (anonymous):

So.... you need to find an equivalent fraction with larger numbers? Or you need a larger fraction?

OpenStudy (alexandervonhumboldt2):

to find which fraction is larger you need to make all denominators the same :)

OpenStudy (electronia):

I failed. I had a time limit.